OpenStudy (anonymous):

What are the discontinuities of the function f(x) = the quantity x squared minus 16 over the quantity 6x minus 24 ? @jim_thompson5910

OpenStudy (anonymous):

bottom can not be 0, which means that 6 x can not equal 24 means that you x = 4 is your discontinuity

jimthompson5910 (jim_thompson5910):

dariobl is correct set the denominator equal to zero and solve 6x - 24 = 0 6x = 24 x = 24/6 x = 4 makes the denominator zero, so this causes the discontinuity

OpenStudy (anonymous):

thanks